# What is the Funniest Number? An investigation of numerical humor

### v1

@@ Introduction
There are a few infamous numbers that are broadly considered, at least in Internet culture, to be funny. Some examples of such numbers are 69 [4], 420 [5], and (531)8008 [6]. These are what we will call mimetically funny numbers, which are associated with or symbolic of a taboo or otherwise humorous topic. The humor of these numbers is not grounded in any sort of numeracy but is instead derived from the direct connection to a concrete, tangible concept. As this study is interested in the former, we take care to differentiate between numerically and mimetically funny cases. One interesting note can be made on these numbers in the Benign Violation Theory framework–numbers themselves are, despite the topic of this work, not generally thought of as funny. The fact that mimetically funny numbers are funny (insofar as they bring to mind a humorous topic) is therefore a violation of the general perception of numbers. Most people do not care enough about numbers for this violation to be anything but benign. Numbers are not supposed to be funny, so the fact that they are funny makes them funny.
@@ Results
Participants showed some difference in their responses for the Zero Context portion compared to the Minimal Context portion: for all responses, and after removing trivial responses. Some participants commented on the difference between the Zero Context and Minimal Context portions of the survey in the optional prompt for explanations and general thoughts at the end of the survey:
@@
To test the PSH, we consider the fraction of responses for the combined survey data (plotted in Figure 3) according to the degree of specificity as defined previously. According to the survey, the funniness of a number is associated with degree of specificity: and . This association is less significant in the Minimal Context data, although the significance substantially increases for both sets of data upon removing responses of 0: . The level of significance is consistent after removing trivial responses. It is clear from Figure 4 that participants were more likely to select numbers in the given range with higher degrees of specificity.

### v2

@@ Introduction
There are a few infamous numbers that are broadly considered, at least in Internet culture, to be funny. Examples of such numbers include 69 [4], 420 [5], and (531)8008 [6]. These are what we will call mimetically funny numbers, which are associated with or symbolic of a taboo or otherwise humorous topic. The humor of these numbers is not grounded in any sort of numeracy but is instead derived from the direct connection to a concrete, tangible concept. As this study is interested in the inherent humor of numbers, we take care to differentiate between numerically and mimetically funny cases. One interesting note can be made on these numbers in the Benign Violation Theory framework–numbers themselves are, despite the topic of this work, not generally thought of as funny. The fact that mimetically funny numbers are funny (insofar as they bring to mind a humorous topic) is therefore a violation of the general perception of numbers. Most people do not care enough about numbers for this violation to be anything but benign. Numbers are not supposed to be funny, so the fact that they are funny makes them funny.
@@ Results
Participants showed some difference in their responses for the Zero Context portion compared to the Minimal Context portion: for all responses, and after removing trivial responses. Some participants commented on the difference between the Zero Context and Minimal Context portions in the optional prompt for explanations and general thoughts at the end of the survey:
@@
To test the PSH, we consider the fraction of responses for the combined survey data (plotted in Figure 3) according to the degree of specificity as defined previously. According to the survey, the funniness of a number is associated with the degree of specificity: and . This association is less significant in the Minimal Context data, although the significance substantially increases for both sets of data upon removing responses of 0: . The level of significance is consistent after removing trivial responses. It is clear from Figure 4 that participants were more likely to select numbers in the given range with higher degrees of specificity.